OpenStudy (anonymous):

(03.01) If the cubic function P(x) includes the points (−4, 0), (0, 0), and (2, 0), which of the following represents this function? x(x − 4)(x + 2) −4(x − 4)(x + 2) x(x + 4)(x − 2) 2(x + 4)(x − 2)

OpenStudy (anonymous):

Need a lot of help with this!!!!!

OpenStudy (anonymous):

Please answer and walk me through it thanks in advance!!!!!!!!!!

OpenStudy (anonymous):

EXPLANATION POINTS!!!!!!!

OpenStudy (anonymous):

Here is a few more I can't stand

OpenStudy (anonymous):

I will show you how to check the first equation, but the rest of the work is down to you. The equation is:\[ y=x(x − 4)(x + 2)\]So we need to check if (−4, 0), (0, 0) and (2, 0) fit this equation. First, try (-4,0):\[-4=0(0-4)(0+2)=0\]Obviously wrong as 4 doesn't equal 0. No need to test the other points as this doesn't lie on the line. Do this for the other 3 and you'll find your answer.

OpenStudy (anonymous):

(03.01) If the polynomial P(x) has roots −2, 1, 1, and 3, which of the following represents the factored form of function P(x)? (x − 1)(x − 1)(x + 2)(x − 3) (x + 1)(x + 1)(x − 2)(x + 3) (x − 1)(x + 1)(x + 2)(x − 3) (x + 1)(x − 2)(x + 3)

OpenStudy (anonymous):

So B?????

OpenStudy (anonymous):

#tom982

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

@fbi2015

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

Is the first one B

OpenStudy (anonymous):

No. Read what I wrote and work it out, don't guess.

OpenStudy (anonymous):

Im not guessing

OpenStudy (anonymous):

I believe 4 and 3 are wrong

OpenStudy (anonymous):

Try again, one of those is the answer.